A college student earned $7200 during summer vacation working as a waiter at a popular restaurant. the student invested part of the money at 9% and the rest at 6%. if the student received a total of $492 in interest at the end of the year, how much was invested at 9%?
Answer by ReadingBoosters(3246)

.09x + .06y = 492
x+y=7200
....................
Substitution
.09(7200-y) + .06y = 492
648 - .09y + .06y = 492
648 - 492 = .09y - .06y
156 = .03y
5200 = y = $ at 6%
7200 - 5200 = 2000 = x = $ at 9%
....................
$5200 at 6%
$2000 at 9%